# Knowledge Adaptation as Posterior Correction

### v1

@@ 3.4 Federated Learning as Posterior Correction
$q_{i}=\arg\min_{q_{i}\in\mathcal{Q}}\,\,\mathbb{E}_{q_{i}}\left[\ell_{i}+\sum_{j=1:t,j\neq i}\hat{\ell}_{i|t}\right]+\mathbb{D}_{\text{KL}}[q_{i}\,\|\,p_{0}].$ (22)
Proceeding in this manner, if the algorithm converges, we recover the joint model both locally and globally, that is, $q_{\text{jnt}}=\hat{q}_{\text{jnt}}=q_{i}$ for all $i$. The PVI update is essentially a scheme to iteratively minimize the gradient mismatch, eventually driving it to 0. The connection ultimately shows that PVI aims for a perfect posterior correction.
There are better alternatives to PVI and they too are connected to posterior correction. A recent work by Swaroop et al. (2025) connects PVI to Federated Alternating Direction Method of Multipliers (ADMM) (Gabay and Mercier, 1976; Glowinski and Marroco, 1975) and its variants for federated deep learning, for example, FedADMM (Gong et al., 2022; Wang et al., 2022; Zhou and Li, 2023) and FedDyn (Acar et al., 2021). They modify the PVI update to include a learning rate $1/\gamma$ in front of the KL term in Eq. 21. This simple change makes it similar to the update of Tseng (1991) and improves convergence. The scheme simply replaces the stale ^$\hat{\ell}_{i|t}^{\text{old}}$ by a moving average $\gamma\ell_{i|t}+(1-\gamma)\ell_{i|t}^{\text{old}}$.
Möllenhoff et al. (2025) further show that ADMM can also be derived as a special case of a more general Bayesian Duality. Essentially, they replace the global update shown in Eq. 19 by the following which employs a learning rate $\alpha=\gamma/(\gamma+t)$,
@@
$\hat{q}_{\text{jnt}}\propto\,\,\left[\prod_{i=1}^{t}q_{i}^{1/t}\right]^{\alpha}\left[p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i|i}\right)\right]^{(1-\alpha)}$ (23)
This schedule is shown to work well for ADMM and Möllenhoff et al. (2025) also show improvements over PVI. In our view, such updates further stabilizes the local update of individual ^$\hat{\ell}_{i|t}$ and improves convergence. This connection also enable us to write posterior correction as optimizing a Lagrangian, therefore using an algorithm similar to the well-known primal-dual methods; interested readers are encouraged to see Möllenhoff et al. (2025). These extensions bring new insight into the existing Bayesian literature (Vehtari et al., 2020; Ashman et al., 2022) which is currently missing a rigorous connection to decade old work in convex duality and distributed optimization. Altogether, with such connections, posterior correction opens a new way to design better adaptive algorithms.
@@ 4.1 A Dual Perspective of The Bayesian Learning Rule (BLR)
Posterior correction is derived by using a dual perspective of the Bayesian Learning Rule (BLR) by Khan and Rue (2023). The perspective is similar to those used in the optimization literature, for example, the closed-circuit of Rockafellar (1967, Fig. 2) which also have parallels in the Bayesian literature, for example, the representer theorem by Kimeldorf and Wahba (1970). Both use ‘dual’ representation of solutions. The representer theorem in Kernel methods generalizes them to a pure functional form by exploiting convex duality (Schölkopf et al., 2001). The dual-perspective of the BLR extends such results to the VL problem.
The form in Eq. 3 represents the solution $q_{t}$ as a product of multiple factors, traditionally called sites in the approximate Bayesian literature (Minka, 2001). The expression is derived by using the expectation parameters $\boldsymbol{\mu}=\mathbb{E}_{q}[\mathbf{T}(\boldsymbol{\theta})]$. The pair $(\boldsymbol{\lambda},\boldsymbol{\mu})$ live in two spaces dual to each other, that is, they are connected by a bijective Legendre transform (Khan and Rue, 2023, Sec 2.2). The natural-gradient are simply gradients in the $\boldsymbol{\mu}$-space which simplifies their computations. The dual form is obtained by simply setting $\widetilde{\nabla}$ of Eq. 1 to 0,
@@
The first expression follows by rewriting the KL, the second equation from the fact that ~$\widetilde{\nabla}\mathcal{H}(q_{t})=-\boldsymbol{\lambda}_{t}$, and the third equation is obtained by applying $\exp(\langle\mathbf{T}(\boldsymbol{\theta}),\cdot\rangle)$ on both sides. For non-constant base measures $h(\boldsymbol{\theta})$, only a small change in the second step is required due to the change in the natural gradient of the entropy term (Khan and Rue, 2023, App. B). We choose to handle it through the site of $\ell_{0}$.
@@
The derivation reveals the underlying dual structure: The left side in the second expression belongs to the $\boldsymbol{\lambda}$ space while the gradients in the right are computed in the dual $\boldsymbol{\mu}$ space. This is akin to the dual structure used in Rockafellar (1967, Fig. 2) and other forms of representer theorems. A formal derivation can be done via a Lagrangian formulation which shows that natural gradients can also be seen as Lagrange multipliers. We skip this discussion as it is not our main focus in this paper, but interested readers may see similar formulations in Khan et al. (2013); Adam et al. (2021); Möllenhoff et al. (2025).
@@
As shown in Table 1, the site functions for Gaussian take a similar form to surrogates obtained by using Taylor’s method. We now derive some of those. We first consider the $q^{\text{iso}}_{t}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mbox{$\mbox{$\mathbf{I}$}$})$ for which we have $\mathbf{T}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\boldsymbol{\mu}_{t}=\mathbb{E}_{q_{t}}[\boldsymbol{\theta}]=\mathbf{m}_{t}$. With this, we get
@@
^$\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})=\langle\boldsymbol{\theta},\nabla_{\mathbf{m}}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}],$ (25)
@@
where we pushed the gradient inside by using the Bonnet’s theorem (Bonnet, 1964; Khan and Rue, 2023). We can also write a similar expression for a $q^{\text{full}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mbox{$\mbox{$\mathbf{S}$}$}_{t}^{-1})$ for which we have $\mathbf{T}(\boldsymbol{\theta})=(\boldsymbol{\theta},\boldsymbol{\theta}\boldsymbol{\theta}^{\top})$. Using this, we get
@@
^$\begin{split}\hat{\ell}^{\text{full}}_{i|t}(\boldsymbol{\theta})&=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle\\ &=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla\ell_{i}-(\nabla^{2}\ell_{i})\mathbf{m}_{t}\right]+\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla^{2}\ell_{i}\right]\boldsymbol{\theta}\\ &=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\mathbb{E}_{q_{t}}\left[\nabla^{2}\ell_{i}\right](\boldsymbol{\theta}-\mathbf{m}_{t})+\text{const.}\end{split}$ (26)
@@
where in the first line we used the expression for the natural gradients given in Khan and Rue (2023, Eq. 11). In the derivations above, we have ignored the constants but these expressions are essentially a result of a $1^{\text{st}}$-order Taylor expansion in the (lifted) $\boldsymbol{\mu}$-space used in the BLR during optimization. For instance, a BLR step uses the following $1^{\text{st}}$-order expansion of $\mathbb{E}_{q}[\ell_{i}]$ for an arbitrary $q$ with the parameter $\boldsymbol{\mu}$:
@@
$\mathbb{E}_{q}[\ell_{i}]\approx\mathbb{E}_{q_{t}}[\ell_{i}]+\langle\boldsymbol{\mu}-\boldsymbol{\mu}_{t},\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle=\mathbb{E}_{q}\left[\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle\right]+\text{const.}$.
Such approximation in the $\boldsymbol{\mu}$-space taken by the BLR yield the $1^{\text{st}}$ and $2^{\text{nd}}$ order expansion in the $\boldsymbol{\theta}$-space as special cases.
@@ 4.2 Posterior Correction as Bayes’ Filter and Natural-Gradient Mismatch
The posterior correction in Eq. 4 is derived using the variational form but the update can also be written in a form similar to Bayes’ rule. Essentially, by using the dual form of $q_{t+1}$, we can directly express it in terms of $q_{t}$ as follows,
$\begin{split}q_{t+1}\propto\left(e^{-\hat{\ell}_{t+1|t+1}}\right)\times q_{t}\times\prod_{i=0}^{t}\left(e^{-\hat{\ell}_{i|t+1}+\hat{\ell}_{i|t}}\right)\end{split}$ (27)
The recursive form is similar to Bayesian filtering but there is an additional third term to account for the interference introduced in the past $\ell_{i}$ due to the inclusion of $\ell_{t+1}$ (a similar update for federated learning is in Eq. 20). Using this, we can also write the divergence to $q_{t+1}$ from $q_{t}$ by simply rearranging to get $\mathbb{E}_{q_{t+1}}[\log(q_{t+1}/q_{t})]$,
$\mathbb{D}_{\text{KL}}[q_{t+1}\,\|\,q_{t}]=\mathbb{E}_{q_{t+1}}\left[-\hat{\ell}_{t+1|t+1}\right]-\sum_{i=0}^{t}\mathbb{E}_{q_{t}}\left[\hat{\ell}_{i|t+1}-\hat{\ell}_{i|t}\right]+A(\boldsymbol{\lambda}_{t})-A(\boldsymbol{\lambda}_{t+1}).$ (28)
@@
This expression can be seen as an extension of the Information Gain (Lindley, 1956). The gain is defined for the exact posterior $p_{t}$ and, when updated to $p_{t+1}$ by using new $\ell_{t+1}$, it is equal to $\mathbb{E}_{p_{t+1}}[-\ell_{t+1}]$. The first term in Eq. 28 is similar but uses the site instead of the loss. The second and third terms are due to the correction. To the best of our knowledge, no such closed-form expression exist to the date to characterize information gain of variational posteriors. For small gains, adaptation should be quick. Large ones may take longer. This is how posterior correction quantifies the feasibility of quick adaptation.
@@
We can use $q_{\text{chk}}$ to boost the training trajectories of the BLR where we simply replace the prior $p_{0}$ by the $q_{\text{chk}}$ and correct the losses $\ell_{i}$ accordingly. This is shown below where we simplify the BLR update of Khan and Rue (2023, Eq. 22) taken at $q_{\text{old}}$,
$\begin{split}q_{\text{new}}&=\arg\min_{\boldsymbol{\mu}}\,\,\boldsymbol{\mu}^{\top}\widetilde{\nabla}\left(\sum_{i=1}^{t}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,p_{0}]\right)+\frac{1}{\rho}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,q_{\text{old}}]\\ &=\arg\min_{\boldsymbol{\mu}}\,\,\sum_{i=1}^{t}\boldsymbol{\mu}^{\top}\widetilde{\nabla}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}]+\frac{1}{\alpha}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,p_{0}^{\alpha}q_{\text{old}}^{1-\alpha}]\\ &=\arg\min_{\boldsymbol{\mu}}\,\,\sum_{i=1}^{t}\boldsymbol{\mu}^{\top}\widetilde{\nabla}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}-\hat{\ell}_{i|\text{chk}}]+\frac{1}{\alpha}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,{q_{\text{chk}}}^{\alpha}q_{\text{old}}^{1-\alpha}].\end{split}$ (30)
The first line is obtained by simply plugging in the VL objective. The second line is simplified by noting that the KL between $q_{\boldsymbol{\mu}}$ and $p_{0}$ contains conjugate term so it can be simply taken out and merged with the last KL term by redefining $\alpha=\rho/(1+\rho)$. An update of this form is in Khan and Rue (2023, Eq. 59) and a derivation for the conjugate prior case is given in Nickl et al. (2023, Eq. 26). The final step is obtained by also noting the fact that ^$\hat{\ell}_{\text{chk}}$ is also a conjugate form, so it can simply be inserted inside the first term.
The derivation above shows the strength of the posterior correction. Not only that it can be applied to trained model, it can also be used to boost the training. The check-point also need not belong to a trained model. As long as we can express them in a dual form, we can use their stored knowledge to boost training of other models. This particular feature makes posterior correction a fundamental mechanism to reuse and repurpose the knowledge learned during model training.

### v3

@@ 3.4.3 Partition Variational Inference as Posterior Correction
$\begin{split}&q_{i}\leftarrow\operatorname*{arg\,min}_{q_{i}\in\mathcal{Q}}\,\,\mathbb{E}_{q_{i}}\left[\ell_{i}+\log s_{i}\right]+\mathbb{D}_{\text{KL}}[q_{i}\,\|\,q_{\text{jnt}}]\\ &s_{i}^{\text{old}}\leftarrow s_{i},\,\,\text{ and }\,\,s_{i}\leftarrow s_{i}^{\text{old}}\frac{q_{i}}{q_{\text{jnt}}}\\ &q_{\text{jnt}}\leftarrow\frac{1}{\mathcal{Z}_{\text{jnt}}}q_{\text{jnt}}\prod_{i=1}^{t}\frac{s_{i}}{s_{i}^{\text{old}}}.\end{split}$ (32)
We can show that, for EF posteriors, these three steps equivalent to Eq. 31 if we set $\rho=1$. For such cases, the sites are simply $s_{i}\propto\exp(-\hat{\ell}_{i})$. A detailed derivation of the equivalence is in Sec. G.4.
The choice $\rho=1$ in PVI may not be appropriate, as also noted by ashman2022partitioned, the algorithm can diverge without additional “damping”. An appropriate range of $\rho$ which ensures convergence is discussed in Ts91. For instance, when using a quadratic regularizer as in Eq. 29, the range $\rho\in(0,2/t)$ ensures convergence under mild assumptions on $\ell_{i}$. Clearly, for $t>2$, this range never includes $\rho=1$. To improve stability, ashman2022partitioned introduced a damping factor $\rho\in(0,1)$ to update $s_{i}$ but this is not used in the update of $q_{i}$. It is unclear whether the resulting procedure converges or not, but it does slow down the update; see further discussion in Sec. G.4. In practice, we do observe cases where PVI can diverge (moellenhoff2025) and recommend a careful use of $\rho$ by using the suggestions of Ts91.
3.4.4 ADMM-Style Extensions of Federated Posterior Correction
It is also possible to use updating schemes for posterior corrections that are similar to Alternating Direction Method of Multipliers (ADMM) (GaMe76; GlMa75). A major difference between ADMM and AMA is in the former’s update of $\boldsymbol{\theta}_{\text{jnt}}$ where an additional sum over $\boldsymbol{\theta}_{i}$ is included, as shown below,
$\boldsymbol{\theta}_{\text{jnt}}\leftarrow\alpha\sum_{i=1}^{K}\mbox{$\mbox{$\mathbf{v}$}$}_{i}+(1-\alpha)\sum_{i=1}^{K}\boldsymbol{\theta}_{i},$ (33)
where $\alpha=1/(1+\rho K)$. The federated posterior correction can easily be modified to include such updating, for instance, we can replace the update of $q_{\text{jnt}}$ in Eq. 26 by the following:
@@ 3.4.4 ADMM-Style Extensions of Federated Posterior Correction
$q_{\text{jnt}}\leftarrow\frac{1}{\mathcal{Z}_{\text{jnt}}}\left[p_{0}\prod_{i=1}^{t}\exp\left(-\hat{\ell}_{i}\right)\right]^{\alpha}\left[\prod_{i=1}^{t}q_{i}\right]^{1-\alpha}.$ (34)
This update can also be interpreted as a BLR step at $\prod_{i}q_{i}$ with learning rate $\alpha$. A similar extension of the BLR is recently proposed by moellenhoff2025 who also use the dual form of the VB solution, referred to as the Bayesian Duality structure. They are the first to derive both AMA and ADMM as special cases of a federated Bayesian algorithm based on such a dual structure. They also propose a modification of PVI, as well as a new variant of the algorithm by SwKh25, and present their ADMM variants, such as, FedADMM (gong2022fedadmm; wang2022fedadmm; ZhLi23) and FedDyn (acar2021feddyn). These extensions bring new insight into the existing Bayesian literature (vehtari2020expectation; ashman2022partitioned) which is currently missing a rigorous connection to decade old work in convex duality and distributed optimization. Altogether, with this type of connections, posterior correction opens a new way to design better adaptive algorithms.
@@ 4.1 A Dual Perspective of The Bayesian Learning Rule (BLR)
Posterior correction is derived by using a dual perspective of the Bayesian Learning Rule (BLR) by khanRue23. The duality is similar to those used in the optimization literature, for example, the closed-circuit of Ro67. Such perspectives do have roots in the Bayesian literature, for example, the representer theorem by kimeldorf1970correspondence was original proposed in a Bayesian context. Unfortunately, these ideas are rarely used in the modern Bayesian machine-learning community, but they remains popular in the kernel-methods community (scholkopf2001generalized). The dual-perspective of the BLR connects them back to Bayes via the variational-Bayesian route.
In this section, we give additional details about the dual form shown in Eq. 3. The form represents the solution $q_{t}$ as a product of multiple factors, traditionally referred to as ‘sites’ in the graphical-model literature. The expression is derived by using the expectation parameters $\boldsymbol{\mu}=\mathbb{E}_{q}[\mathbf{T}(\boldsymbol{\theta})]$. The pair $(\boldsymbol{\lambda},\boldsymbol{\mu})$ live in two spaces dual to each other and connected by a bijective Legendre transform (khanRue23, Sec 2.2). As a consequence, the natural-gradient in the $\boldsymbol{\lambda}$-space are simply gradients in the $\boldsymbol{\mu}$-space, that is, $\widetilde{\nabla}_{\boldsymbol{\lambda}}=\nabla_{\boldsymbol{\mu}}$ (khan2018fast1, Thm. 1). This fact is used in the BLR to simplify the computation of natural gradients.
With this, the dual form in Eq. 3 is obtained by applying $\widetilde{\nabla}$ to Eq. 1 and setting to 0. The calculation is shown below where we rewrite $\mathbb{D}_{\text{KL}}[q_{t}\,\|\,p_{0}]=\mathbb{E}_{q_{t}}[\ell_{0}]-\mathcal{H}(q_{t})$ where $\mathcal{H}(q_{t})$ is the entropy, then get the second equality by using the fact that $\widetilde{\nabla}\mathcal{H}(q_{t})=-\boldsymbol{\lambda}_{t}$, and write the third equation by substituting $\boldsymbol{\lambda}_{t}$ in $q_{t}\propto\exp(\langle\mathbf{T}(\boldsymbol{\theta}),\boldsymbol{\lambda}_{t}\rangle)$:
@@
The derivation in Eq. 36 reveals the underlying dual structure: The left side in the second expression belongs to the $\boldsymbol{\lambda}$ space while the gradients in the right are computed in the dual $\boldsymbol{\mu}$ space. This is akin to the dual structure used in Ro67 and other forms of representer theorems. A formal derivation can be done via a Lagrangian formulation which shows that natural gradients can also be seen as Lagrange multipliers. We skip this discussion as it is not our main focus in this paper, but interested readers may see similar formulations in KhAr13; adam2021dual; moellenhoff2025. The extension to distributions with non-constant base measure is discussed in App. B.
The dual form also holds during learning when using the BLR to optimize the variational objective. The BLR uses the following moving average with a learning rate $\rho$, which in a similar fashion gives rise to a dual form that holds during training,
$\boldsymbol{\lambda}_{t}\leftarrow(1-\rho)\boldsymbol{\lambda}_{t}+\rho\sum_{i=0}^{t}\widetilde{\nabla}\mathbb{E}_{q_{t}}[-\ell_{i}]\qquad\implies\qquad q_{t}\leftarrow\frac{1}{\mathcal{Z}_{t}}q_{t}^{1-\rho}\prod_{i=0}^{t}\exp\left(-\rho\hat{\ell}_{i|t}\right).$ (36)
A logistic regression example is included in Sec. A.3. One can go even further and write a dual form by ‘unrolling’ all the updates to an older iteration, say $j$ step back in the past. This is similar to the ‘lazy’ form used in online learning (hoeven2018many) and was originally proposed by khan2017conjugate; see also khanRue23.
@@
We will now derive the explicit form of the site function for Gaussian distributions shown in Table 1. For such cases, the sites take a similar form to surrogates obtained by using Taylor’s method. We first consider the $q^{\text{iso}}_{t}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mbox{$\mbox{$\mathbf{I}$}$})$ for which we have $\mathbf{T}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\boldsymbol{\mu}_{t}=\mathbb{E}_{q_{t}}[\boldsymbol{\theta}]=\mathbf{m}_{t}$. Using this, we can write
@@
$\hat{\ell}^{\text{iso}}_{i|t}(\boldsymbol{\theta})=\langle\boldsymbol{\theta},\nabla_{\mathbf{m}}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}],$ (37)
@@
where in the last expression we push the gradient inside the expectation by using the Bonnet’s theorem (bonnet1964transformations; khanRue23). We can also write a similar expression for a $q^{\text{full}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m}_{t},\mbox{$\mbox{$\mathbf{S}$}$}_{t}^{-1})$ for which we have $\mathbf{T}(\boldsymbol{\theta})=(\boldsymbol{\theta},\boldsymbol{\theta}\boldsymbol{\theta}^{\top})$. Using this, we get
@@
$\begin{split}\hat{\ell}^{\text{full}}_{i|t}(\boldsymbol{\theta})&=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle\\ &=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla\ell_{i}-(\nabla^{2}\ell_{i})\mathbf{m}_{t}\right]+\mbox{$\frac{1}{2}$}\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}\left[\nabla^{2}\ell_{i}\right]\boldsymbol{\theta}\\ &=\boldsymbol{\theta}^{\top}\mathbb{E}_{q_{t}}[\nabla\ell_{i}]+\mbox{$\frac{1}{2}$}(\boldsymbol{\theta}-\mathbf{m}_{t})^{\top}\mathbb{E}_{q_{t}}\left[\nabla^{2}\ell_{i}\right](\boldsymbol{\theta}-\mathbf{m}_{t})+\text{const.}\end{split}$ (38)
@@
where in the first line we make use of the expression given in khanRue23. In the derivations above, we have ignored the constants but these expressions are essentially a result of the first-order Taylor expansion in the (lifted) $\boldsymbol{\mu}$-space used in the BLR during optimization. For instance, denoting $L_{i}(\boldsymbol{\mu})=\mathbb{E}_{q}[\ell_{i}]$, we apply a mirror-descent update which linearizes $L_{i}$ in the $\boldsymbol{\mu}$-space,
@@
$\begin{split}\boldsymbol{\lambda}_{t}&\leftarrow\arg\min_{\boldsymbol{\mu}}\sum_{i=0}^{t}\left[L_{i}(\boldsymbol{\mu}_{t})+\langle\boldsymbol{\mu}-\boldsymbol{\mu}_{t},\nabla_{\boldsymbol{\mu}}L_{i}(\boldsymbol{\mu}_{t})\rangle\right]-\mathcal{H}(q_{t})+\frac{1}{\rho}\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]\\ &=\arg\min_{\boldsymbol{\mu}}\sum_{i=0}^{t}\mathbb{E}_{q}\langle\boldsymbol{\mu},\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle+\frac{1}{\rho}\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}^{1-\rho}],\end{split}$ (39)
where we ignored the constants and absorbed the entropy term in the KL term. This form is originally proposed in khan2017conjugate. Information geometry based explanations are given in khan2025information.
@@ 4.2 Posterior Correction as Bayes’ Filter and Natural-Gradient Mismatch
The posterior correction in Eq. 4 is derived using the variational form but the update can also be written in a form similar to Bayes’ rule. Essentially, we apply the dual form given in Eq. 3 to Eq. 4 which gives the following form for $q_{t+1}$ in terms of $q_{t}$,
$\begin{split}q_{t+1}=\frac{\mathcal{Z}_{t}}{\mathcal{Z}_{t+1}}\left(e^{-\hat{\ell}_{t+1|t+1}}\right)\times q_{t}\times\prod_{i=0}^{t}\left(e^{-\left(\hat{\ell}_{i|t+1}-\hat{\ell}_{i|t}\right)}\right)\end{split}$ (40)
The recursive form is similar to Bayesian filtering but there is an additional third term to account for the interference introduced in the past $\ell_{i}$ due to the inclusion of $\ell_{t+1}$ (a similar update for federated learning is in Eq. 27). Using this, we can also write the divergence to $q_{t+1}$ from $q_{t}$ by simply rearranging Eq. 40 to write $\mathbb{E}_{q_{t+1}}[\log(q_{t+1}/q_{t})]$ as follows,
$\mathbb{D}_{\text{KL}}[q_{t+1}\,\|\,q_{t}]=\mathbb{E}_{q_{t+1}}\left[-\hat{\ell}_{t+1|t+1}\right]-\sum_{i=0}^{t}\mathbb{E}_{q_{t}}\left[\hat{\ell}_{i|t+1}-\hat{\ell}_{i|t}\right]+\log\frac{\mathcal{Z}_{t}}{\mathcal{Z}_{t+1}}.$ (41)
@@
This expression can be seen as an extension of the Information Gain (lindley1956measure), which is defined as $\mathbb{D}_{\text{KL}}[p_{t+1}\,\|\,p_{t}]=\mathbb{E}_{p_{t+1}}[-\ell_{t+1}]+\log(\mathcal{Z}_{t}/\mathcal{Z}_{t+1})$ where $p_{t}$ and $p_{t+1}$ are the old and new posteriors respectively. The first term in Eq. 41 is similar but uses the site instead of the loss. The second term is additional and arises due to the correction. To the best of our knowledge, no such closed-form expression exist to the date to characterize information gain of variational posteriors. For smaller gains, adaptation should be quicker. This is how posterior correction quantifies the feasibility of quick adaptation.
@@
We can use $q_{\text{chk}}$ to boost the training trajectories of the BLR where we simply replace the prior $p_{0}$ by the $q_{\text{chk}}$ and correct the losses $\ell_{i}$ accordingly. This is shown below where we simplify the BLR update of khanRue23 taken at $q_{\text{old}}$,
$\begin{split}q_{\text{new}}&=\operatorname*{arg\,min}_{\boldsymbol{\mu}}\,\,\boldsymbol{\mu}^{\top}\widetilde{\nabla}\left(\sum_{i=1}^{t}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,p_{0}]\right)+\frac{1}{\rho}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,q_{\text{old}}]\\ &=\operatorname*{arg\,min}_{\boldsymbol{\mu}}\,\,\sum_{i=1}^{t}\boldsymbol{\mu}^{\top}\widetilde{\nabla}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}]+\frac{1}{\alpha}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,p_{0}^{\alpha}q_{\text{old}}^{1-\alpha}]\\ &=\operatorname*{arg\,min}_{\boldsymbol{\mu}}\,\,\sum_{i=1}^{t}\boldsymbol{\mu}^{\top}\widetilde{\nabla}\mathbb{E}_{q_{\boldsymbol{\mu}}}[\ell_{i}-\hat{\ell}_{i|\text{chk}}]+\frac{1}{\alpha}\mathbb{D}_{\text{KL}}[q_{\boldsymbol{\mu}}\,\|\,{q_{\text{chk}}}^{\alpha}q_{\text{old}}^{1-\alpha}].\end{split}$ (43)
The first line is obtained by simply plugging in the VL objective. The second line is simplified by noting that the KL between $q_{\boldsymbol{\mu}}$ and $p_{0}$ contains conjugate term so it can be simply taken out and merged with the last KL term by redefining $\alpha=\rho/(1+\rho)$. An update of this form is in khanRue23 and a derivation for the conjugate prior case is given in nickl2023memory. The final step is obtained by also noting the fact that $\hat{\ell}_{i|\text{chk}}$ is also a conjugate form, so it can simply be inserted inside the first term.
The derivation above shows the strength of the posterior correction. Not only that it can be applied to trained model, it can also be used to boost the training. The check-point also need not belong to a trained model. As long as we can express them in a dual form, we can use their stored knowledge to boost training of other models. This particular feature makes posterior correction a powerful mechanism to reuse and repurpose the knowledge learned during model training.